OpenStudy (anonymous):

Find x if the average of 20, 20, 19, 13, and x is 16. A. 6 B. 8 C. 9 D. 10

OpenStudy (anonymous):

b

OpenStudy (anonymous):

b is correct

OpenStudy (anonymous):

(20+20+19+13+x)/5=16 72+x=80 x=8,b

OpenStudy (anonymous):

How did you get 72+x=80 ????

OpenStudy (anonymous):

add 20 20 19 13 gives 72 multiply 16 by 5 gives 80

OpenStudy (anonymous):

I see how to do it now. Thank you @MAYUR2345 and @imyint .

OpenStudy (anonymous):

Can you guys help me with another one please?

OpenStudy (anonymous):

While attending a film festival you decide that there are 13 movies that you are interested in seeing. However, you only have time to see 7 movies. In how many different ways can you watch 7 of the 13 movies? A. 8,648,640 B. 13 C. 665,280 D. 4,324,320

OpenStudy (anonymous):

a calculate 13P7